wiki:UC20190305

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 3/05/19

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), John Wisniewski (OU), Joanne Hughes (Seattle U), Anne Verbiscer (UVa), Mary Beth Kaiser (JHU), John Bally (CU), Russet McMillan (APO), Moire Prescott (NMSU), Misty Bentz (GSU), Chip Kobulnicky (UWy), Eric Bellm (UW), Ben Williams (UW)


User feedback and comments from institutional representatives:

  • Georgia State - nothing to report
  • Colorado - nothing to report
  • NAPG - no report
  • JHU - nothing to report
  • NMSU - nothing to report
  • Seattle - nothing to report
  • Oklahoma - wanted to pass along to the APO day staff that their willingness to help the OU observational techniques class do cals during the daytime is much appreciated
  • Washington - nothing to report
  • Wyoming - nothing to report
  • UVa - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 2/04/19 - 2/28/19

0) Overview

Winter weather has not been particularly good; the site has seen lots of clouds and humidity but little additional snow accumulation. Overall minimal problems, which went along with minimal on-sky time. However, the DIS problems did dominate engineering’s time.

1) Telescope

The telescope had no problems aside from occasional guider issues that generally cleared with restarting or INITs.

2) Instruments

DIS - The DIS blue camera is now in the clean room and we found a couple of shorted wires for the preamp board. The preamp board was checked and is OK but we won't have an idea of the status of the CCD until it is reassembled and connected up at the telescope. If the electronics look OK warm we will then pump and cool the camera in about a week. Only then will we know the complete state of the CCD. We should have some more updates on the Users’ Committee telecon. The DIS red camera failed to give images but during troubleshooting suddenly recovered and is working as of this date. Note that we did not purposely do anything to correct the problem.

TripleSpec is operational.

Agile is operational.

Echelle is operational. We started initial commissioning of the new Finger Lakes slit viewer camera but ran into a problem where 5-10% of the images were shifted. Troubleshooting is in progress.

NIC-FPS was operational. The ICC died 20 minutes ago and we will start troubleshooting on March 1st and have update for the Users’ Committee telecon.

ARCTIC is operational.

3) ARCSAT

Operational without guide camera. Some upgrades to the computer and communications were attempted during engineering time. Unfortunately the older hardware cannot support some of these upgrades so we reverted many of the items. We are still looking at why the guider camera is not transferring images correctly through new communication hardware.


Additional telescope and instrument discussion:

The main issue has still been with DIS. DIS Red failed and then came back online after the last servicing; it recovered without any intervention from the engineering staff. The problem was nothing drastically serious (but we don’t know what it was, which means it could reoccur at any time). The symptom was that on DIS Red biases were normal if not exposed to light, but if exposed to light the chip wouldn’t respond or would have strange readout streaks. DIS Blue was having problems with noise after the last servicing, then after the engineering staff went in and put it back together they found further problems with its readout. They eliminated all possibilities if the problem were outside dewar. When they opened up the dewar they found some electrical problems (shorted wires from preamp board), which may have been caused when adding desiccant into the dewar (which was intended to mitigate the issues with contamination). We are waiting for someone to come back from leave to deal with a wire that needs to be resoldered and terminated. Then we will reassemble everything and hook it up electrically while still warm to see if the signal levels seen before are corrected. If so they will bring it back downstairs, pump it down and cool it (this will take roughly 5 days). If we put everything back together and the signal levels are still bad then likely the chip got damaged and we will explore putting in a spare. We are hoping to know more later this week. DIS Red is still available for use, and scattered light on the red side is still “very good.”

We've had some problems with the new slit viewer camera for echelle, as well as shutter problems on the spectrograph; we are currently exploring options. This is indicative of ongoing problems with guide cameras/slit viewers in general. We don't plan to do any drastic engineering that would reduce the instrument's availability because it is currently taking up the slack due to the aforementioned DIS issues. The replacement slit viewer needs a driver. The old one isn’t getting worse so we can still use it for the time being. The shutter problem is with the echelle spectrograph, but replacing it is not trivial; shutters that large are quite expensive. The issue only affects observations of bright sources, including cal lamps. The problem is that it is sticking open, which manifests itself as streaks in the vertical direction as charge clocks across the chip. An observer would see streaks above or below the Ar lines in cal lamp exposures. The frame would look blurry if looking at a continuum source (similar to the inter-order light problem for targets with bright emission lines, which would get smeared into adjacent orders).

Did some work on ARCSAT - see the Telescope and Instrument Report above for more details. Upgrading it was not as successful as we would have liked because of old hardware. We are exploring options, but we are still a long way from having constructive recommendations for upgrade plans.

The NICFPS ICC died. We had hoped that we had a machine on site that could work when we transferred the hard disk over, but this was not successful, suggesting that there must be something unique in the communication chain in the existing hardware. We ordered a new replacement power supply, which should arrive next week. We are still looking at ways to make this hardware work but it is slow going since the ICC was delivered by the instrument builders and not something that was put together locally. Furthermore, the observatory is still operating with limited computing support pending a new hire. Bill reached out to the software developer who wrote all the code for NICFPS, but he has moved on and is no longer part of the CU department, and has not responded to Bill. John Bally confirmed that the original team that built NICFPS has dissipated.


Q1 3.5-m scheduling:

We have requests for all remaining outstanding time for March except Mar 31, which we have been saving for overflow from April.


Q1 ARCSAT scheduling:

There is some open time the last week of Mar. We will entertain Q2 schedule requests after the 3.5m schedule is finished. Maybe target Mar 15 for a call for proposals?


Q2 3.5-m scheduling:

We are just starting to build the Q2 schedule. Bravo to everyone for getting their institutional requests in on time. There are a couple of conflicts so far and there may be more. There were lots of restrictive and/or confusing requests; Russet sent out emails seeking clarification.


NA2 Baffle Project:

Nancy reported that John Wilson has requested that the ARCTIC web site be updated to reflect the true FOV of the instrument. Otherwise no news on this project.


KOSMOS Update:

The KOSMOS Critical Design Review telecon was held in late February. On that call there was a lot of discussion concerning the instrument cart, mounting plate, slits, internal calibration lamps, etc. Those who were requested to provide feedback are asked to do so by end of this week. The consensus after the CDR telecon was that the UW team will proceed with the design efforts for enabling observations of internal calibration lamps. The slit viewer is also proceeding, although we still need to quantify how well matched it will be to the limiting magnitude of the science instrument. Construction of the new (longer hold time) dewar is moving along; they have already cut the metal. The control software/operations/electronics are still progressing but it is still TBD as far as what it will look like. A graduate student at UW is working on the electronics side, Conor is working on the software side. No show-stoppers were identified at the CDR stage. We expect to commission the instrument at APO in June, and we are setting aside some time in the Q2 schedule for this purpose. We are interested in having people suggest some science targets to assess the instrument’s capabilities and performance during commissioning.


Response to Call for White Papers for 2020 Decadal Survey

There is a telecon scheduled tomorrow to discuss this.


Discussion of 25th Anniversary / Strategic Planning meeting in 2019

Nancy sent out the survey to poll people in terms of preferred dates and locations. The deadline for responses has passed; she is now in the process of scheduling the meeting (date, location) based on that feedback, so stay tuned for more details soon.


ACTION ITEMS:

Open action items from previous meetings:

  • Nancy will send out an update as soon as anything else is known about the DIS blue camera. STATUS: CLOSED.
  • Misty Bentz and John Bally will upload their NICFPS data reduction scripts to the wiki. STATUS: CLOSED/OPEN. Misty has done this; John will do it.
  • Ben Williams will ask Anjum whether she had any Agile data reduction documentation she could upload to the wiki. STATUS: OPEN. Ben got Anjum's new email address and contacted her about this. Although she has been pretty inactive lately, she said that she is interested in supplying something but she hasn’t sent anything yet. Ben will send her a reminder.
  • Nancy will disseminate the TEWG final report. STATUS: OPEN.
  • Nancy will organize a telecon to discuss the Astro2020 white paper. STATUS: OPEN.
  • Users Committee reps should discuss the 2019 strategic planning meeting with their departments, identify possible attendees, and have them complete the survey before 2/18/19. STATUS: CLOSED.

New action items from this meeting:

  • Nancy will coordinate Astro2020 science white paper submission based on telecon tomorrow.
  • Nancy will disseminate information concerning the date/location of the Strategic Planning meeting in June 2019.

Next meeting:

The next Users Committee telecon will be on Monday April 1, 2019, at 10:00 am MDT.

Note: See TracWiki for help on using the wiki.